OpenStudy (anonymous):

what is the answer in x squared - 2x - 3 in using polynomial

OpenStudy (anonymous):

\[x^2 - 2x - 3 = 0 ?\]

OpenStudy (anonymous):

\[x^2 - 3x + x - 3=0\] \[x(x-3) + (x-3)=0\] \[(x+1)(x-3)= 0\] x=-1, x=3

OpenStudy (anonymous):

katrose : ( x-1 ) x+3 )